# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 4.2 A Roadmap for an Aspiring Market Designer
Step 9. Forge Strategic Partnerships with Insiders: While you may have started this endeavor as a critical “outsider,” you cannot complete it while remaining in this role. To influence policy and institute change, you need to collaborate with insiders. Forming these partnerships will transition you from an outsider to an insider, enabling you to transform your innovations into concrete policies and potentially paving the way for more elaborate interventions.
@@
One effective method I’ve found for building trust is providing pro-bono assistance. Attempting to sell your services without this gesture may not reassure stakeholders about your intentions. [^35] Charging for your services is acceptable, or even expected, if insiders recruit you as a consultant for this endeavor. However, as an outsider seeking to change an institution, offering pro-bono support emphasizes your sincerity, highlighting a key difference between commissioned and aspired market design.
@@
Step 10. Tinker with the Design, Incorporating Additional Input from Stakeholders if Necessary: Once a partnership is established with authorities, you are no longer bound by the constraints that directly influenced both your research questions and proposed design, as outlined in Steps 1-9. After all, you are no longer an outsider. Make the most of this added flexibility.
@@
Note that this final step fundamentally differs from the earlier ones. Steps 1-9 can be viewed as efforts aimed at convincing decision makers that you are a “worthy” partner in this endeavor—the persuasion phase of minimalist market design. Consequently, many elements of the framework are deliberately “conservative.” This conservatism helps our work to be seen as an interdisciplinary collaboration rather than an intrusive approach, significantly enhancing its reception among experts in other fields and decision makers (see Section 14.1).
However, once a certain level of trust is established between the relevant parties, the approach in this final step need not be as conservative. Similarly, the role of theory can be toned down at this stage. In this sense, there is potentially a more robust role in Step 10 for alternative methodologies, including computational, experimental, and empirical economics. Thus, the open-ended tinkering step in minimalist market design can be seen as the stage where you transition to a “commissioned” role, guiding the design with greater flexibility. [^36] It is tempting to associate Step 10 of my framework with the “tinkering” phase of policy design in Esther Duflo’s influential Richard T. Ely Lecture “The Economist as Plumber” (Duflo, 2017). While this association is natural, there are some important distinctions between the roles of the tinkering phase in the two frameworks. In Duflo’s framework, the role of the tinkering phase is making tweaks to prescribed policy to correct any issues that may arise in field implementation. Such issues can emerge due to the gap between theoretical analysis and the actual problem. Thus, the tinkering phase is particularly valuable in settings where the primary role of theory is building intuition. In my framework, in contrast, there should not be a mismatch between theory and practice at a level that could lead to any serious issues. The role of the tinkering phase is further improving the institution with the additional flexibility gained by the design economist at this stage.
@@
Having outlined a guide to minimalist market design, in the next part of the monograph, Sections 5 through 12, I will present several applications that exemplify the holistic interaction between research and policy within this framework.
5 Residential Assignment
@@
Before the mid-1990s, two related problems had been studied in the literature. The first, the housing markets model by Shapley and Scarf (1974), represents a unit-demand exchange economy where each individual owns a single indivisible good—a “house”—and has preferences over all houses. The second, the house allocation model by Hylland and Zeckhauser (1979), represents a unit-demand fair allocation problem where a set of “collectively owned” houses is allocated among individuals based on their preferences. The primary distinction between the two models lies in the property rights.
@@ 5.2 Housing Markets and the Core
But for me, the most amazing thing about this model is the following procedure—the Gale’s Top Trading Cycles (G–TTC) algorithm—which generates the unique outcome of a housing market. This algorithm was famously invented by David Gale and suggested to Herbert Scarf during one of Scarf’s early presentations of the housing markets model (Scarf, 2009).
Consider any housing market. The key technical insight driving the G–TTC algorithm is this: there is always at least one cycle of individuals—possibly consisting of a single person—where each person prefers the endowment house of the next person in the cycle as their first choice. Consequently, all individuals in any cycle can receive their first choices by pooling their initial endowments and reassigning them among themselves.
So, why not form such a cycle in the first round of an algorithm, awarding all individuals in each cycle their first choices, and keep repeating this process with the individuals remaining in the market and their endowment houses? What a brilliant idea!
But why does such a cycle exist in every housing market? Let’s explore this through an example.
@@
Next, consider Banu—the person who owns Alp’s first-choice house. Her first choice is either her own house, forming a one-person cycle, or it’s Alp’s house, forming a two-person cycle with Alp. If neither, her first choice must be Cora’s house, since no one else remains.
Now, Cora’s top choice must be the house of one of the three. If it’s her own, she forms a one-person cycle. If it’s Alp’s, she forms a two-person cycle with Alp. And if it’s Banu’s, she forms a three-person cycle with Banu and Alp.
@@
Amidst all this beauty, the only downside to this elegant model was that, at the time, it didn’t seem to have any real-world application. Little did I know that my efforts, which began at the SITE conference, would help change that years later. Through a “discovery-invention cycle” driven by minimalist market design—starting with Shapley and Scarf’s housing market model and G–TTC, and continuing with the key innovations by Abdulkadiroğlu and Sönmez (1999)—the ideas behind Gale’s brilliant algorithm would eventually lead to saving thousands of lives every year through living-donor organ exchanges worldwide (see Section 7.2).
@@ 5.3 House Allocation and Simple/Random Serial Dictatorship
The key difference between these two models lies in property rights: while the first model represents a private-ownership economy, the second represents a collective-ownership economy.
A straightforward solution for the house allocation model is to mimic the natural dynamics of a queue. Students are first placed in a sequence based on criteria such as seniority, grade point average (GPA), or a random lottery. Then, housing units are allocated one at a time according to the students’ positions in the queue and their preferences. The first student in line receives their top choice, the next student gets their top choice from the remaining units, and so on.
When this procedure is used with a fixed sequence of students, it is known as a simple-serial dictatorship (SSD) (Satterthwaite and Sonnenschein, 1981). When it is applied to a randomly ordered sequence of students, uniformly distributed, it is called a random serial dictatorship (RSD) (Abdulkadiroğlu and Sönmez, 1998).
SSD and RSD mechanisms perform well regarding their efficiency and incentive-compatibility properties. SSD is Pareto efficient, while RSD is ex-post Pareto efficient, meaning its outcome—a probability distribution among deterministic outcomes—gives positive weight only to Pareto efficient outcomes. Both mechanisms are also strategy-proof.
@@ 5.4 House Allocation with Existing Tenants
Graduate housing is typically awarded for several years, and during the annual allocation process, students who already occupy units from the previous year are entitled to keep them if they choose. Thus, in this version of the problem, there are “existing tenants” who are earmarked their “occupied” units, “newcomers” without an earmarked unit, and “vacant” units.
Motivated by this more complex setting, Abdulkadiroğlu and Sönmez (1999) models and studies this variation, calling it house allocation with existing tenants. This model generalizes both the housing markets model by Shapley and Scarf (1974), where all units are occupied and assigned to their current occupants, and the house allocation model by Hylland and Zeckhauser (1979), where no units are earmarked, and all are available for allocation.
@@ 5.5 Simple/Random Serial Dictatorship with Squatting Rights
Even though the house allocation model does not capture all aspects of the practical on-campus housing problem, system operators often adapt solutions from this simpler setting to address their needs in more complex scenarios. Indeed, this approach—relying on adaptations from simpler models to solve more challenging problems—is not limited to authorities overseeing real-life applications; it is also a common method among design economists. In this vein, many institutions, including Carnegie Mellon, Duke, Harvard, Northwestern, and Pennsylvania, have implemented a straightforward modification of the SSD (or RSD) to accommodate the added complexity of existing tenants.
@@ 5.6 You Request My House–I Get Your Turn: A Minimalist Redesign
Although SSD and RSD possess attractive efficiency and incentive-compatibility properties in standard house allocation, SSD (or RSD) with Squatting Rights sacrifices these qualities when adapted to accommodate existing tenants. Since existing tenants risk receiving a less desirable unit than their current one if they relinquish it to join the applicant pool, these mechanisms generally do not incentivize them to give up their earmarked units. This can reduce potential gains from trade, ultimately compromising efficiency.
@@
To prevent this issue, whenever Alp’s first choice is the earmarked unit of another existing tenant, we move that tenant to the top of the queue, right before Alp, and proceed with the adjusted order. In this case, the queue becomes Banu, Alp, Cora, and Diya.
@@
With the new queue, no matter what Cora’s first choice is, awarding her this unit cannot compromise individual rationality. Depending on Cora’s choice, either a “cycle” forms or a “chain” that starts with the vacant house:
@@
Regardless of the specifics of the problem, this simple adjustment to the dynamics of SSD—where an individual who has not yet received their assignment is promoted to the top of the queue once another individual shows interest in their earmarked unit—always generates an individually rational outcome through similar cycles or chains. The chain may start either with vacant units or with units whose occupants have already left the queue for a more preferred unit.
This minimalist redesign, first introduced by Abdulkadiroğlu and Sönmez (1999) and known as the You Request My House–I Get Your Turn (YRMH–IGYT) algorithm, resolves the shortcomings of SSD with Squatting Rights when used for house allocation problems with existing tenants as a direct mechanism.
@@
Another key feature of the YRMH–IGYT mechanism is its ability to accommodate any deterministic or probabilistic priority order of individuals, similar to a serial dictatorship. This feature provides system operators with an additional policy lever. For example, in on-campus housing, the system can assign higher priority to upperclassmen, students with higher grades, or students from disadvantaged backgrounds.
As we progress through this monograph, we will see that the YRMH–IGYT algorithm—one of my first key inventions—plays a crucial role in multiple “discovery-invention cycles” discussed throughout the text. To fully appreciate its significance, let’s explore this procedure in greater depth.
@@
Under these property rights, the YRMH–IGYT algorithm enhances the basic SSD by integrating potential welfare gains through two types of trades. First, similar to the G–TTC algorithm, it facilitates the exchange of earmarked units among existing tenants through “cycles” of individuals. Second, it enables the trade of a leading spot in the queue—whether held by an existing tenant or a newcomer—for an earmarked unit of an existing tenant. This latter trade creates a “chain” initiated by the queue leader at various stages, who then receives either a vacant unit or an unassigned unit vacated by an earlier processed existing tenant.

### v4

@@ 4.2 A Roadmap for an Aspiring Market Designer
Step 9. Forge Strategic Partnerships with Insiders.
While you may have started this endeavor as a critical “outsider,” you cannot complete it while remaining in this role. To influence policy and institute change, you need to collaborate with insiders. Forming these partnerships will transition you from an outsider to an insider, enabling you to transform your innovations into concrete policies and potentially paving the way for more elaborate interventions.
@@
One effective method I’ve found for building trust is providing pro-bono assistance. Attempting to sell your services without this gesture may not reassure stakeholders about your intentions. [^41] Charging for your services is common, and often expected, if insiders recruit you as a consultant for this endeavor. However, as an outsider seeking to change an institution, offering pro-bono support emphasizes your sincerity, highlighting a key difference between commissioned and aspired market design.
@@
Step 10. Tinker with the Design, Incorporating Additional Input from Stakeholders if Necessary.
Once a partnership is established with authorities, you are no longer bound by the constraints that directly influenced both your research questions and proposed design, as outlined in Steps 1–9. After all, you are no longer an outsider. Make the most of this added flexibility.
@@
Note that this final step fundamentally differs from the earlier ones. Steps 1–9 can be viewed as efforts aimed at convincing decision-makers that you are a “worthy” partner in this endeavor—the persuasion phase of minimalist market design. Consequently, many elements of the framework are deliberately “conservative.” This conservatism helps our work to be seen as an interdisciplinary collaboration rather than an intrusive approach, significantly enhancing its reception among experts in other fields and decision-makers (see Section 15.1).
However, once a certain level of trust is established between the relevant parties, the approach in this final step need not be as conservative. Similarly, the role of theory can be toned down at this stage. In this sense, there is potentially a more robust role in Step 10 for alternative methodologies, including computational, experimental, and empirical economics. Thus, the open-ended tinkering step in minimalist market design can be seen as the stage where you transition to a “commissioned” role, guiding the design with greater flexibility. [^42] It is tempting to associate Step 10 of my framework with the “tinkering” phase of policy design in Esther Duflo’s influential Richard T. Ely Lecture “The Economist as Plumber” (Duflo, 2017). While this association is natural, there are some important distinctions between the roles of the tinkering phase in the two frameworks. In Duflo’s framework, the role of the tinkering phase is making tweaks to prescribed policy to correct any issues that may arise in field implementation. Such issues can emerge due to the gap between theoretical analysis and the actual problem. Thus, the tinkering phase is particularly valuable in settings where the primary role of theory is building intuition. In my framework, in contrast, there should not be a mismatch between theory and practice at a level that could lead to any serious issues. The role of the tinkering phase is further improving the institution with the additional flexibility gained by the economic designer at this stage.
@@
Having outlined a guide to minimalist market design, in the next part of the monograph, Sections 5 through 12, I will present several applications that exemplify the holistic interaction between research and policy within this framework. [^44] Although this monograph is not intended as a textbook, Sections 5 through 12 (or a subset thereof) are highly structured and contain sufficient formalism and examples to support an advanced undergraduate or graduate course on matching markets or market design. These sections can be supplemented with Sönmez and Ünver (2025a, b) to increase the mathematical and technical depth of the material.
5 Housing Allocation
@@
Before the mid-1990s, two related problems had been studied. The first, the housing markets model of Shapley and Scarf (1974), is an exchange economy in which each individual is endowed with a single indivisible good—a “house”—and has preferences over all houses. Individuals may either keep their houses or trade them if they wish. The second, the house allocation model of Hylland and Zeckhauser (1979), is an allocation problem in which a central planner distributes houses among individuals according to their preferences. The key distinction between the two lies in property rights: in the first, houses are privately endowed, whereas in the second, individuals may have equal claims or be priority-listed according to various criteria.
@@ 5.2 Housing Markets and the Core
But for me, the most amazing thing about this model is the following procedure, which generates the unique core outcome of a housing market. This algorithm was famously invented by David Gale and suggested to Herbert Scarf during one of Scarf’s early presentations of the housing markets model (Scarf, 2009).
Gale’s Top Trading Cycles (G–TTC) Algorithm.
The key technical insight of the G–TTC algorithm is this: there is always at least one cycle of individuals—possibly consisting of a single person—where each person prefers the endowment house of the next person in the cycle as their first choice. Consequently, all individuals in any cycle can receive their first choices by pooling their initial endowments and reassigning them among themselves. So, why not form such a cycle in the first round of an algorithm, awarding all individuals in each cycle their first choices, and keep repeating this process with the individuals remaining in the market and their endowment houses? What a brilliant idea!
But why does such a cycle exist in every housing market? We explore this through an example, illustrated in Figure 5.484848Throughout the monograph, individuals are denoted by the first letter of their names in the figures (e.g., Alp is denoted as A, Banu as B, Cora as C, etc.).
Figure 5: Existence of a cycle in Gale’s Top Trading Cycles Algorithm (Example 1). Label the first agent as A. If A’s top choice is their own house (Case a), A immediately forms a one-person cycle. Otherwise, A’s top choice is another house, whose owner we label B (Cases b and c). If B’s top choice belongs to someone already encountered (either A or B), a cycle arises. In particular, if B prefers their own house (Case b1), there is a one-person cycle, whereas if B prefers A’s house (Case b2), then B and A form a two-person cycle. If instead B’s top choice is C’s house (Case c), then C’s top choice must belong to someone already encountered, since no one else remains. In this case, if C prefers their own house (Case c1), there is a one-person cycle; if C prefers B’s house (Case c2), then C and B form a two-person cycle; and if C prefers A’s house (Case c3), then C, B, and A form a three-person cycle. Thus, in every scenario, a cycle emerges, denoted with bold arrows.
@@
Next, consider Banu—the owner of Alp’s first-choice house. Her first choice is either her own house, forming a one-person cycle, or it’s Alp’s house, forming a two-person cycle with Alp. If neither, her first choice must be Cora’s house, since no one else remains.
Now, Cora’s top choice must be the house of one of the three. If it’s her own, she forms a one-person cycle. If it’s Banu’s, she forms a two-person cycle with Banu. And if it’s Alp’s, she forms a three-person cycle with Alp and Banu.
@@
Amidst all this beauty, the only downside to this elegant model was that, at the time, it didn’t seem to have any real-world application. Little did I know that my efforts, which began at the SITE conference, would help change that years later.
Through a “discovery–invention cycle” driven by minimalist market design—starting with Shapley and Scarf’s housing market model and G–TTC, and continuing with the key innovations of Abdulkadiroğlu and Sönmez (1999)—the ideas behind Gale’s brilliant algorithm would eventually lead to saving thousands of lives every year through living-donor organ exchanges worldwide (see Section 7.2 and Figure 37).
@@ 5.3 House Allocation and Simple/Random Serial Dictatorship
The key difference between these two models lies in their property-rights structures: the first represents a private-ownership economy, whereas in the latter individuals may either be priority listed according to various criteria or have equal claims.
A straightforward solution for the house allocation model is to mimic the natural dynamics of a queue. Students are placed in a sequence based on priority criteria such as seniority, grade point average (GPA), or, in the case of equal claims, a random lottery. Housing units are then allocated one at a time according to the students’ positions in the queue and their preferences: the first student receives their top choice, the next receives their top remaining choice, and so on.
Serial Dictatorship.
This procedure, when applied with a fixed sequence of students, was first studied by Hylland and Zeckhauser (1979) and later termed a serial dictatorship by Satterthwaite and Sonnenschein (1981). In subsequent work, it has sometimes been referred to as a simple serial dictatorship (SSD), which is the terminology we will adopt throughout this monograph. When the sequence of students is drawn uniformly at random, the mechanism is known as random serial dictatorship (RSD), a term introduced by Abdulkadiroğlu and Sönmez (1998).
The SSD and RSD mechanisms perform well with respect to their efficiency and incentive-compatibility properties: SSD is Pareto efficient, RSD is ex post Pareto efficient—its outcome being a probability distribution over deterministic outcomes that assigns positive weight only to Pareto-efficient ones—and both mechanisms are strategy-proof (Svensson, 1994). [^49] Svensson (1994) focuses on SSD, referring to its outcomes as queue allocations.
@@ 5.4 House Allocation with Existing Tenants
Graduate housing is typically awarded for several years, and in the annual allocation process, students who already occupy units from the previous year are entitled to keep them if they choose. Thus, in this version of the problem, there are “existing tenants” with their earmarked “occupied” units, “newcomers” without an earmarked unit, and “vacant” units.
Motivated by this more complex setting, Abdulkadiroğlu and Sönmez (1999) model and study this variation, calling it house allocation with existing tenants. This model generalizes both the housing markets model of Shapley and Scarf (1974), in which all units are earmarked, and the house allocation model of Hylland and Zeckhauser (1979), in which no units are earmarked and all are available for allocation.
@@ 5.5 Simple/Random Serial Dictatorship with Squatting Rights
Even though the house allocation model does not capture all aspects of the practical on-campus housing problem, system operators often adapt solutions from this simpler setting to address their needs in more complex scenarios. Indeed, this approach—relying on adaptations from simpler models to solve more challenging problems—is not limited to authorities overseeing real-life applications; it is also a common method among economic designers. In this vein, many institutions, including Carnegie Mellon, Duke, Harvard, Northwestern, and Pennsylvania, have implemented a straightforward modification of the SSD (or RSD) to accommodate the added complexity of existing tenants.
@@ 5.6 You Request My House–I Get Your Turn: A Minimalist Redesign
Although SSD and RSD feature attractive efficiency and incentive-compatibility properties in standard house allocation, SSD (or RSD) with Squatting Rights loses these qualities when adapted to accommodate existing tenants. Because existing tenants risk receiving a less desirable unit than their current one if they relinquish it to join the applicant pool, these mechanisms generally do not incentivize them to give up their earmarked units. This reduces potential gains from trade and ultimately compromises efficiency.
@@
To prevent this issue, whenever Alp’s first choice is a unit earmarked for another existing tenant— as in this case with Banu— we move that tenant to the top of the queue, right before Alp, and proceed with the adjusted order. The queue then becomes Banu, Alp, Cora, and Diya.
@@
With the new queue, no matter what Cora’s first choice is, awarding her this unit cannot compromise individual rationality. Depending on Cora’s choice, either a “cycle” forms or a “chain” terminating at the vacant house:
@@
Regardless of the specifics of the problem, this simple adjustment to the dynamics of SSD—where an individual who has not yet received an assignment is promoted to the top of the queue once another individual requests their occupied unit—always yields an individually rational outcome through cycles or chains. Such a chain may terminate either at a vacant unit or at a unit whose occupant has already left the queue for a more preferred one.
This resulting minimalist redesign, first introduced in Abdulkadiroğlu and Sönmez (1999), is called the You Request My House–I Get Your Turn (YRMH–IGYT) algorithm.
To further illustrate how cycles and chains arise in this algorithm, we next present its execution with a more structured example.
Example 3.
There are five existing tenants Alp, Banu, Cora, Diya, and Ezra, each endowed with an occupied house $H_{A}$, $H_{B}$, $H_{C}$, $H_{D}$, and $H_{E}$, respectively. There is also a newcomer, Frank, with no endowment, and a vacant house $V$ that is not tied to any individual. All individuals are priority-ordered in a queue in reverse alphabetical order of their names. Preferences of individuals over houses are given as follows:
$\begin{array}[]{llll}\mbox{Alp}:&H_{B}-V-H_{D}-H_{C}-H_{A}-H_{E}\qquad\quad&\mbox{Diya}:&H_{C}-H_{A}-H_{D}-V-H_{B}-H_{E}\\ \mbox{Banu}:&V-H_{A}-H_{E}-H_{C}-H_{D}-H_{B}&\mbox{Ezra}:&V-H_{B}-H_{A}-H_{D}-H_{E}-H_{C}\\ \mbox{Cora}:&H_{B}-H_{E}-H_{D}-V-H_{C}-H_{A}&\mbox{Frank}:&H_{C}-H_{A}-H_{B}-H_{E}-V-H_{D}\end{array}$
We execute YRMH–IGYT, beginning with Frank in Step 1, the first individual in the queue. His top choice is $H_{C}$, the endowment of Cora who is still in the line. Cora is therefore promoted to the front, and the adjusted queue becomes: Cora, Frank, Ezra, Diya, Banu, Alp.
In Step 2, Cora’s top choice is $H_{B}$, the endowment of Banu, who is also still in the line. Banu moves to the front, yielding an adjusted queue: Banu, Cora, Frank, Ezra, Diya, Alp.
In Step 3, Banu’s top choice is the vacant house $V$. This creates a chain: Banu receives $V$, Cora takes $H_{B}$ which has just been vacated by Banu, and Frank takes $H_{C}$ which has just been vacated by Cora. Once this chain exits, the remaining individuals are lined up as follows: Ezra, Diya, Alp.
In Step 4, Ezra’s top choice among the remaining houses is $H_{A}$, the endowment of Alp who is still in the line. Alp is promoted to the top, yielding the further adjusted queue: Alp, Ezra, Diya.
In Step 5, Alp’s top choice among the remaining houses is $H_{D}$, the endowment of Diya who is still in the line. Diya is moved ahead, resulting in the adjusted queue: Diya, Alp, Ezra.
In Step 6, Diya’s top choice among the remaining houses is $H_{A}$, which generates a 2-person cycle with Alp: Diya takes $H_{A}$ and Alp takes $H_{D}$. Both then leave the process.
In Step 7, the only person that remains is Ezra and the only house that remains is his endowment $H_{E}$. His top choice is therefore his own house, forming a 1-person cycle, and he leaves with $H_{E}$.
See Figure 6 for the execution of the YRMH–IGYT algorithm.
Figure 6: Execution of YRMH–IGYT for Example 3. Arrows to individuals indicate a request for their endowment house, an arrow to vacant house V indicates a request for it, and arrows to the front of the queue indicate a priority upgrade to the front of the queue for the next step. Step 3 features a chain, Step 6 a 2-person cycle, and Step 7 a single-person cycle, all depicted with bold arrows.
The execution illustrates how chains and cycles naturally emerge under YRMH–IGYT, leading to a structure that helps address the shortcomings of SSD with Squatting Rights. The next proposition highlights those properties of YRMH–IGYT that are especially useful in contrasting the mechanism with its real-life counterpart.
@@
Another key feature of the YRMH–IGYT mechanism is its flexibility to accommodate any deterministic or probabilistic priority order of individuals, similar to a serial dictatorship. This feature provides system operators with an additional policy lever. For example, in on-campus housing, the system can assign higher priority to upperclassmen, students with higher grades, or students from disadvantaged backgrounds.
As we progress through this monograph, we see that the YRMH–IGYT algorithm—one of my early key inventions—plays a crucial role in multiple “discovery–invention cycles” discussed throughout the text (see Figures 36 and 37). To fully appreciate its significance, let’s explore this procedure in greater depth.
@@
Under these property rights, the YRMH–IGYT algorithm enhances the basic SSD by integrating potential welfare gains through two types of trades. First, similar to the G–TTC algorithm, it facilitates the exchange of earmarked units among existing tenants through “cycles” of individuals. Second, it enables the trade of a leading spot in the queue—whether held by an existing tenant or a newcomer—for an earmarked unit of an existing tenant. This latter trade creates a “chain” initiated by the queue leader at various stages, who then receives either a vacant unit or an unassigned unit vacated by an existing tenant who has already exited the procedure with their assignment.